OpenStudy (bananas1234):

Find the zeros of the polynomial function. f(x) = x^2 + 4x – 60

OpenStudy (zale101):

Set the function equal to zero. Factor and find the x-intercepts.

OpenStudy (zale101):

Those x values determines where the graph crosses the x-axis.

OpenStudy (bananas1234):

ok i thought so : D

OpenStudy (zale101):

Can you show me how did you do it? If you have found the zeroes of the polynomials, then your question has been answered.

OpenStudy (bananas1234):

i wish it was that simple ha ha @Nnesha

OpenStudy (bananas1234):

ok -10 and 6?

OpenStudy (zale101):

Yes, you are correct \(x=\Large \frac{-4\pm\sqrt{4^2-4(1)(-60)}}{2(1)}=\) x=6, and -10

OpenStudy (bananas1234):

thank you